I have tried to find the solution by addition for this equation but my answer comes up wrong. The answer key lets me know that it is dependent but I don't see how. Can you help me?
Here is the problem:
2x-4y=11
-4x+8y=-22

Answer by stanbon(75887)

2x-4y=11
-4x+8y=-22
-----
Comment:: Dependent means the two equation are really the same.
----
Multiply thru the top equation by -2 and you get:
-4x + 8y = -22
That is the same as the bottom equation.
One is simply another way of expressing the other.
They are dependent.
Cheers,
Stan H.